OpenStudy (anonymous):

@AnswerMyQuestions @iGreen FIND THE SUM OR DIFFRENCE 15/16-7/13 A.13/32 B.23/32 C.3/4

OpenStudy (igreen):

What's the least common denominator between 16 and 13?

OpenStudy (anonymous):

3

OpenStudy (igreen):

The least common denominator can't be smaller than the two numbers.

OpenStudy (igreen):

Well the least common multiple here will be, \(15 \times 13 = 208\). So we change the fractions into denominators of \(208\): \(\dfrac{195}{208} - \dfrac{105}{208}\) \(\dfrac{195-105}{208}\) Can you subtract 195 - 105?

OpenStudy (anonymous):

90

OpenStudy (igreen):

Hmm. I think there's something wrong with your question.

OpenStudy (anonymous):

YES THE FRACTION 7/13 IS SPOSE TO BE 7/32

OpenStudy (igreen):

Um..oh..

OpenStudy (igreen):

So the lcd is actually 32.

OpenStudy (anonymous):

OKA SORRY ABOUT THAT

OpenStudy (igreen):

We change 15/16 to a denominator of 32. \(\dfrac{15}{16} \rightarrow \dfrac{15 \times 2}{16 \times 2} \rightarrow \dfrac{30}{32}\) \(\dfrac{30}{32} - \dfrac{7}{32} \rightarrow \dfrac{30-7}{32}\) Can you subtract 30 - 7?

OpenStudy (anonymous):

23

OpenStudy (igreen):

Yep, so what's our answer going to be?

OpenStudy (anonymous):

B